OpenStudy (anonymous):

find the radius of the circle defined by the equation; 3x^2+3y^+42y+186=0 who can help?

OpenStudy (anonymous):

*3y^2

OpenStudy (anonymous):

so?

OpenStudy (anonymous):

dividing with 3 we get x^2+y^2+14y+62=0 hence we can write it as (x-0)^2 +(y+7)^2 +13=0 comparing with standard form answer is sqrt of 13

OpenStudy (anonymous):

This needs a nifty trick known as "Completing the square"

OpenStudy (anonymous):

But yeah, before that, better make sure that the x^2 and y^2 only have coefficients of 1.

OpenStudy (anonymous):

Wait a minute, there's something wrong with this (x-0)^2 +(y+7)^2 +13=0

OpenStudy (anonymous):

I don't think this has any solutions at all... no circle to speak of

OpenStudy (anonymous):

It gives you \[x^2+(y+7)^2=-13\]

OpenStudy (mertsj):

@PeterPan add the 49 to both sides.

OpenStudy (anonymous):

It's been done, @Mertsj That's why 62 became 13

OpenStudy (mertsj):

So there is either a typo or no such circle exists.

OpenStudy (anonymous):

\[x^2+y^2+14y+62=0\]\[x^2+y^2+14y\color{red}{+49}+62=0\color{red}{+49}\]\[x^2+(y+7)^2+62=49\]\[x^2+(y+7)^2+62\color{red}{-62}=49\color{red}{-62}\]\[x^2+(y+7)^2=\color{green}{-13}\]

OpenStudy (anonymous):

Must be a typo @sellweezy96 Please recheck :)

OpenStudy (anonymous):

3x^2+3y^2+42y+186=0

OpenStudy (mertsj):

Could it possibly be -186?

OpenStudy (anonymous):

3x^2+3y^2+42x+42y+186=0 @Mertsj

OpenStudy (anonymous):

Okay, so there's a 42x after all :D

OpenStudy (mertsj):

Well that's a whole different thing.

OpenStudy (anonymous):

First, as @irshadno1 did so shall you do ^.^ Divide everything by 3. We don't want coefficients of x^2 which are not 1.

OpenStudy (anonymous):

x^2+Y^2+14x+14y+62=0 ?

OpenStudy (anonymous):

Good. Now complete the square. x^2 + 14x + y^2 + 14y + 62

OpenStudy (anonymous):

divide it by which

OpenStudy (anonymous):

Here's an example of completing the square \[\large z^2+6z \] See that 6z? Its coefficient is 6. Get half of that, and then square it. Half of 6 is 3, and the square of 3 is 9. Add that. \[\large z^2+6z+9\color{red}{-9}\] \[\large (z+3)^2\color{red}{-9}\]

OpenStudy (anonymous):

So, back to your circle \[\large \color{blue}{x^2+14x}+\color{green}{y^2+14y}+62=0\] You want to complete the square of the blue (x) part?

OpenStudy (anonymous):

=(x+7)+49 ?

OpenStudy (anonymous):

Almost... 14x --> what's half of 14?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

and the square of 7?

OpenStudy (anonymous):

49

OpenStudy (anonymous):

Okay, good :) So you either add and subtract 49 to the left side OR you add 49 on BOTH sides... let's go the other way... \[\large \color{blue}{x^2+14x}\color{red}{+49}+\color{green}{y^2+14y}+62=0\color{red}{+49}\]\[\large \color{violet}{(x+7)^2}+\color{green}{y^2+14y}+62=\color{red}{49}\]

OpenStudy (anonymous):

Now do the same to the green (y) part.